OpenStudy (anonymous):

Approximate the real number solutions to the polynomial function f(x) = x3 – 2x2 – 16x + 32 .

OpenStudy (y2o2):

Use grouping : x³-2x²-16x+32 = x²(x-2) -16(x-2) = (x-2)(x²-16) = 0 so x = 2 or x²=16 [ x = ±4 ]